There is a 50% chance of rain on Saturday and a 50% chance of rain on Sunday. What is the probability of 2 sunny days this weekend?
Answer by stanbon(48516)

There is a 50% chance of rain on Saturday and a 50% chance of rain on Sunday. What is the probability of 2 sunny days this weekend?
--------
The events are independent, so
P(sunny and sunny) = P(sunny)*P(sunny) = 0.5*0.5 = 0.25
===========================================================
Cheers,
Stan H.